What is the amplitude of this function? 2.9 m 4 You are correct. Your receipt no. is 163-4005 Previous Tries What is the period of this function? 1.125s Submit Answer Incorrect. Tries 3/5 Previous Tries What is the frequency (in Hz) of this function? 0.89Hz Submit Answer Incorrect. Tries 3/5 Previous Tries 0 一-1 What is the angular frequency (in rad/s) of this function? 5.585rad/s Incorrect. Tries 1/5 Previous Tries Submit Answer -4 2024 time t [s]
Answer #1

For Given function, Period of function (T) is 2.25 seconds.

Frequency = 1/T = 1/2.25 = 0.44 Hz

And, Angular velocity = 2 (pi) f = 2(3.14)(0.44) = 2.79 rad/s